A researcher reports that the average salary of College Deans is more than Php 63,000. A sample of 35 College Deans has a mean salary of Php 65,700. At a = 0.01, test the claim that the CollegeDeans earn more than Php 63, 000 a month. The standard deviation of the population is Php 5,250. 


1
Expert's answer
2021-06-08T04:28:54-0400

"\mu=63000 \\

n= 35 \\

\bar{x}=65700 \\

\sigma=5250 \\

H_0: \mu ≤ 63000 \\

H_1: \mu > 63000"

The level of significance is "\alpha=0.01"

"z _{crit}=2.326"

The test statistic:

"z=\frac{\bar{x}-\mu}{\sigma / \sqrt{n}} \\

= \frac{65700-63000}{5250/ \sqrt{32}} \\

= 3.043 \\

Z=3.043>Z_{crit}=2.326"

So, we reject H0 and accept H1.

There is evidence to support the claim that the monthly salary of the college dean is more than P63,000